# InvZW: Invariant Feature Learning via Noise-Adversarial Training for Robust Image Zero-Watermarking

### v1

@@ IV-A Dataset and Feature Training
For invariant feature learning and analysis (Section IV-B), we used a subset of 40,000 images from the MSCOCO dataset, which contains a diverse range of real-world scenes. To assess generalization, we evaluated the trained models on 5,000 MSCOCO images, as well as 5,000 images each from the Imagenette and Flickr30k datasets. These datasets introduce unseen variations in content and background, helping to validate the stability of the learned representations under domain shifts. Additionally, high-resolution robustness was tested using 900 images from the DIV2K dataset.
For linear evaluation (Section IV-C), we used 10,000 labeled images from the CIFAR-100 dataset to train a downstream classifier on frozen features, and evaluated it on a separate test set of 3,000 images. CIFAR-100 provides a well-established benchmark for assessing semantic discriminability of features across object categories.
The feature extractor was trained using the Adam optimizer with separate learning rates for each component: $\eta_{\mathrm{FE}}=0.0005$ for the feature extractor, $\eta_{D}=0.001$ for the discriminator, and $\eta_{R}=0.0001$ for the reconstructor. As shown in Figure 5, the training losses of all three components converge stably over time: the discriminator and reconstructor losses decrease steadily, while the feature extractor’s adversarial loss increases as it successfully fools the discriminator, indicating effective minimax optimization and stable training dynamics.
@@
Since the proposed framework learns and generalizes invariance through input distortions, we apply a range of geometric and photometric transformations during training and testing. Table I summarizes the distortion strategies used. Notably, during training, only geometric distortions (e.g., rotation, translation, shear) are applied, while at test time, a broader spectrum of distortions—including blurring, compression, and noise—is introduced. This design enables us to evaluate the generalization capacity of the learned invariant features: although trained only with geometric distortions, the model demonstrates invariance to a wide variety of unseen image processing attacks such as salt-and-pepper noise and JPEG compression.
TABLE I: Summary of Distortions in Training and Testing.
@@
IV-B Analysis of the Proposed Invariant Feature
@@
Table II reports the cosine similarity and PSNR under common image processing distortions. Despite being trained only with geometric augmentations, the model demonstrates strong generalization to unseen distortions. For instance, cosine similarity remains consistently high under blur and JPEG compression ($\geq$0.93 across all datasets), indicating minimal impact on the feature representation. Gaussian noise and solarization result in modest drops in similarity, while salt-and-pepper noise introduces the largest variation, particularly on MSCOCO (0.92). Notably, Flickr30k and ImageNet show higher resilience to these pixel-level attacks, with cosine similarity values reaching up to 0.96.
The PSNR scores corroborate the feature analysis. MSCOCO, ImageNet, and Flickr30k achieve high reconstruction fidelity under most distortions, with blur and JPEG yielding PSNR values above 30 dB. In contrast, DIV2K—due to its higher resolution and finer detail—shows more degradation under extreme perturbations, particularly salt-and-pepper noise (27.3 dB) and solarization (26.0 dB). These results confirm that the learned features are both distortion-invariant and semantically expressive across multiple content domains.
TABLE II: Feature Cosine similarity and Image Reconstruction PSNR under common image processing distortions.
Dataset Cosine Similarity (Feature Invariance) $\uparrow$ PSNR (dB) $\uparrow$
@@
MSCOCO 0.97 0.96 0.96 0.92 0.93 32.5 31.2 30.9 28.8 29.5
ImageNet 0.95 0.94 0.93 0.96 0.92 33.9 30.8 31.5 29.2 30.3
Flickr30k 0.98 0.96 0.95 0.96 0.95 32.2 31.8 29.8 26.5 30.2
DIV2K 0.95 0.93 0.93 0.91 0.93 29.5 31.0 28.7 27.3 26.0
Table III presents the results under geometric transformations. Rotation shows the highest feature invariance across all datasets (cosine similarity $\geq$0.95), followed by width and height shifts. Shear and zoom introduce slightly more disruption, with cosine similarity values dropping to 0.91–0.94 for ImageNet and DIV2K. Among geometric distortions, zoom exhibits the most pronounced impact on feature consistency, likely due to scale-dependent feature extraction challenges.
@@
TABLE III: Feature Cosine similarity and Image Reconstruction PSNR under geometric transformations.
Dataset Cosine Similarity (Feature Invariance) $\uparrow$ PSNR (dB) $\uparrow$
@@
MSCOCO 0.97 0.96 0.96 0.94 0.95 33.5 31.1 31.6 32.3 29.6
ImageNet 0.95 0.94 0.94 0.91 0.93 31.9 33.5 29.2 31.8 29.1
Flickr30k 0.97 0.95 0.94 0.94 0.92 30.2 33.7 31.7 30.2 28.6
DIV2K 0.96 0.92 0.93 0.92 0.94 31.5 32.1 30.8 32.8 29.6
Overall, the results demonstrate that the proposed framework effectively learns invariant features with strong generalization to unseen distortions, including both common image processing effects and geometric spatial transformations. To illustrate that the extracted features preserve not only invariance but also rich semantic information, we validate the reconstruction results on the testing data, and a few representative examples are shown in Figure 6.
@@
IV-C Comparative Analysis of the Proposed Invariant Feature
To evaluate the robustness and effectiveness of the proposed invariant feature learning framework, we compare it with five representative self-supervised learning (SSL) baselines: SimCLR [17], BYOL [18], Barlow Twins [25], VICReg [26], and VIBCREG [27]. These methods were selected for their strong performance across SSL benchmarks and for covering diverse learning paradigms—including contrastive learning, redundancy reduction, and variance-invariance-covariance regularization. We assess the learned representations along two key axes: (1) feature invariance, i.e., the stability of feature embeddings under distortions, and (2) semantic descriptiveness, i.e., how well the features support classification when linearly probed. These two criteria together offer a comprehensive measure of robustness and utility for downstream watermarking tasks.
IV-C1 Invariant Feature Similarity
@@
Table IV reports the results across four datasets. The proposed method consistently outperforms the SSL baselines across all distortions and datasets. For instance, under Gaussian blur and JPEG compression, our model achieves cosine similarities up to 0.98, demonstrating high resistance to low-level degradations. Even under challenging conditions like salt & pepper noise and hue shifts, our model maintains higher feature consistency than all compared methods. These results validate the effectiveness of our adversarial invariant learning objective in producing distortion-invariant feature embeddings.
TABLE IV: Quantitative Comparison of Feature Invariance (Cosine Similarity) across SSL Methods Under Diverse Distortions on Multiple Datasets.
@@
IV-C2 Linear Evaluation for Semantic Discriminability
@@
IV-D Robust Zero Watermarking
This subsection presents an evaluation of the proposed learning-based zero-watermarking scheme using the invariant features extracted in Module 1. We organize the analysis into three parts. First, we visualize the training dynamics of the reference signature $C$, including t-SNE projections before and after optimization, and monitor the convergence behavior of the watermark loss over time. Next, we quantify the robustness of the watermarking system across a wide range of image distortions by evaluating bit accuracy under varying transformation intensities. Finally, we compare our method against existing state-of-the-art zero-watermarking and other deep watermarking approaches, highlighting the performance gains achieved by our framework in terms of robustness and stability. Together, these experiments establish the practicality and generalizability of our proposed method. We evaluated the proposed method using binary watermark messages of varying lengths and report results using a 30-bit watermark to align with common practice in the literature and ensure fair comparison with prior work.
IV-D1 Visualization of Reference Signature Learning
@@
As shown in Fig. 9, we apply t-distributed Stochastic Neighbor Embedding (t-SNE) to project the high-dimensional reference vectors $C\in\mathbb{R}^{k\times d}$ into two dimensions, both before and after training. Each point represents the projection of one bit vector $C_{i}$ associated with a distorted image. Before training, the projected vectors appear scattered without clear structure, reflecting the random initialization. After optimization, the vectors form distinct and compact clusters in the embedding space. This emergent separation indicates that the learning process successfully aligns the reference signature with the distortion-invariant feature space. Notably, the improved spatial organization mirrors the expected behavior of our theoretical formulation, in which $C$ is optimized to serve as a directional codebook that maximizes the margin between binary watermark bits. The optimization implicitly identifies a hyperplane such that $\tilde{F}\cdot C_{i}\geq 0.5$ or $<0.5$, depending on the bit label, enabling robust watermark recovery through simple dot-product thresholding.
@@
IV-D2 Distortion Tolerance Range
@@
For JPEG compression, BER consistently decreases as the quality factor improves, confirming strong robustness to compression artifacts. Gaussian blur exhibits a gradual BER increase with higher blur radii, while cropping and hue adjustment maintain relatively low BER values even at high distortion intensities. This indicates that the proposed method effectively preserves watermark integrity under spatial and color-shift perturbations. Noise-based distortions—especially Gaussian noise and salt & pepper noise—introduce relatively more significant degradation, with BER rising sharply under higher noise levels. Nonetheless, the absolute BER values remain low across all cases, ranging from 0.05% to a maximum of 8%, highlighting the overall robustness of the proposed method.
IV-D3 Comparison with Existing Watermarking Methods
To evaluate the robustness of the proposed watermarking framework, we compare it against several representative models across three categories: (1) deep learning-based watermarking schemes that embed information into the image, (2) feature extractor-based watermarking methods that leverage pre-trained SSL models, and (3) zero-watermarking approaches with deep models. Each comparison focuses on bit-level accuracy or BER under various common distortions.
For a fair comparison, all methods were evaluated using a fixed watermark length of 30 bits per image. Additionally, watermarking embedding was controlled to ensure consistency across methods. In particular, for embedded watermarking schemes, watermarking were applied such that the resulting marked images maintained a PSNR of approximately 40 dB. In contrast, the proposed method and other zero-watermarking baselines do not alter the original image, preserving its fidelity by design.
TABLE V: Bit Accuracy (%) Comparison with Watermarking Models that Embed Information.
@@ IV-D3 Comparison with Existing Watermarking Methods
Table V presents bit accuracy results under a wide range of distortions. We compare our method with HIDDEN [28], a widely used encoder-decoder watermarking scheme, and [29], which uses adversarial training to improve robustness. Our model consistently achieves the highest accuracy across most transformations, particularly under color-related distortions such as hue and saturation adjustment, where alternative methods degrade significantly. Moreover, our model shows stronger robustness under severe noise (e.g., Gaussian 0.10, Salt & Pepper 0.15), outperforming others by large margins. The average accuracy improvement of our model demonstrates superior generalization across both geometric and photometric distortions.
TABLE VI: Bit Accuracy (%) Comparison with Pretrained Feature-Based Watermarking Models.
@@
Table VI compares our framework with two recent methods that utilize pre-trained SSL models for watermark retrieval: Vukotić et al. [7], which uses convolutional networks, and Fernández et al. [8], which employs DINO features. Although these approaches demonstrate strong performance under identity-preserving transformations, they often struggle with heavy attacks and geometric alterations such as rotation or crop. In contrast, our model maintains perfect or near-perfect bit accuracy across all comparison distortions. Notably, under non-differentiable transformations like screenshots, our method preserves 98% bit accuracy, further validating the resilience conferred by our invariant feature learning.
TABLE VII: BER (%) Comparison across Zero-Watermarking Models.
@@
JPEG Compression (30) 0.0267 2.0314
JPEG Compression (70) 0.0169 0.5748
@@
Table VII reports BER comparisons with the recent deep zero-watermarking method from [2]. Across different distortion types—including compression, filtering, and geometric transformations—our method demonstrates competitive or superior robustness. In particular, under geometric distortions such as rotation and scaling, and under common image processing operations like Gaussian filtering, our framework achieves consistently lower BER, with improvements up to 50% in some cases (e.g., scaling at factor 2.0 and Gaussian filter $3\times 3$). While our method shows slightly higher BER under extreme JPEG compression, it maintains strong robustness across a broader range of transformations. These results underscore the generalizability and robustness of the learned invariant features for zero-watermarking under diverse perturbations.
These results collectively demonstrate the superiority of our approach across different methodological paradigms. Unlike existing approaches that rely on embedded information or static feature extractors, our method learns an adaptive and robust signature in an invariant feature space, offering strong performance under a broad range of real-world transformations.
V Conclusion

### v2

@@ 4.1 Dataset and Feature Training
For invariant feature learning and analysis (Section 4.2), we used a subset of 40,000 images from the MSCOCO dataset, which contains a diverse range of real-world scenes. To assess generalization, we evaluated the trained models on 5,000 MSCOCO images, as well as 5,000 images each from the Imagenette and Flickr30k datasets. These datasets introduce unseen variations in content and background, helping to validate the stability of the learned representations under domain shifts. Additionally, high-resolution robustness was tested using 900 images from the DIV2K dataset.
For linear evaluation (Section 4.3), we used 10,000 labeled images from the CIFAR-100 dataset to train a downstream classifier on frozen features, and evaluated it on a separate test set of 3,000 images. CIFAR-100 provides a well-established benchmark for assessing semantic discriminability of features across object categories.
The feature extractor was trained using the Adam optimizer with separate learning rates for each component: $\eta_{\mathrm{FE}}=0.0005$ for the feature extractor, $\eta_{D}=0.001$ for the discriminator, and $\eta_{R}=0.0001$ for the reconstructor. As shown in Figure 5, the training loss curves indicate stable and consistent convergence across all components of our model. The Discriminator and Reconstructor losses decrease progressively, reflecting improved capability in distinguishing invariant features and reconstructing original content. Meanwhile, the Feature Extractor shows a steady increase in performance, suggesting it is learning more robust and invariant representations over time. The smooth convergence of these loss curves highlights the effectiveness of our training strategy and suggests that the model is becoming more accurate and generalizable.
@@
Since the proposed framework learns and generalizes invariance through input distortions, we apply a range of geometric and photometric transformations during training and testing. Table 1 summarizes the distortion strategies used. Notably, during training, only geometric distortions (e.g., rotation, translation, shear) are applied, while at test time, a broader spectrum of distortions, including blurring, compression, and noise, is introduced. To ensure that image rotation does not result in content loss, we apply a combined rotation and resizing operation that preserves the full image content within the frame.This design enables us to evaluate the generalization capacity of the learned invariant features: although trained only with geometric distortions, the model demonstrates invariance to a wide variety of unseen image processing attacks such as salt-and-pepper noise and JPEG compression.
Table 1: Summary of Distortions in Training and Testing.
@@
4.2 Analysis of the Proposed Invariant Feature
@@
Table 2 reports the cosine similarity and PSNR under common image processing distortions. Despite being trained only with geometric augmentations, the model demonstrates strong generalization to unseen distortions. For instance, cosine similarity remains consistently high under blur and JPEG compression ($\geq$0.93 across all datasets), indicating minimal impact on the feature representation. Gaussian noise and solarization result in modest drops in similarity, while salt-and-pepper noise introduces the largest variation, particularly on MSCOCO (0.92). Notably, Flickr30k and ImageNet show higher resilience to these pixel-level attacks, with cosine similarity values reaching up to 0.96.
The PSNR scores corroborate the feature analysis. MSCOCO, ImageNet, and Flickr30k achieve high reconstruction fidelity under most distortions, with blur and JPEG yielding PSNR values above 30 dB. In contrast, DIV2K, due to its higher resolution and finer detail, shows more degradation under extreme perturbations, particularly salt-and-pepper noise (27.3 dB) and solarization (26.0 dB). These results confirm that the learned features are both distortion-invariant and semantically expressive across multiple content domains.
Table 2: Feature cosine similarity and image reconstruction PSNR under common distortions (mean $\pm$ std over 5 runs).
Dataset Cosine Similarity $\uparrow$ PSNR (dB) $\uparrow$
@@
MSCOCO $0.97\pm 0.014$ $0.96\pm 0.024$ $0.96\pm 0.032$ $0.92\pm 0.037$ $0.93\pm 0.021$ $32.5\pm 0.41$ $31.2\pm 0.68$ $30.9\pm 0.83$ $28.8\pm 1.12$ $29.5\pm 0.52$
ImageNet $0.95\pm 0.031$ $0.94\pm 0.038$ $0.93\pm 0.041$ $0.96\pm 0.019$ $0.92\pm 0.036$ $33.9\pm 0.57$ $30.8\pm 1.21$ $31.5\pm 0.92$ $29.2\pm 1.48$ $30.3\pm 1.05$
Flickr30k $0.98\pm 0.011$ $0.96\pm 0.026$ $0.95\pm 0.043$ $0.96\pm 0.031$ $0.95\pm 0.022$ $32.2\pm 0.93$ $31.8\pm 0.74$ $29.8\pm 1.96$ $26.5\pm 1.78$ $30.2\pm 0.88$
DIV2K $0.95\pm 0.028$ $0.93\pm 0.036$ $0.93\pm 0.039$ $0.91\pm 0.044$ $0.93\pm 0.033$ $29.5\pm 1.12$ $31.0\pm 0.91$ $28.7\pm 1.47$ $27.3\pm 1.18$ $26.0\pm 1.32$
Table 3 presents the results under geometric transformations. Rotation shows the highest feature invariance across all datasets (cosine similarity $\geq$0.95), followed by width and height shifts. Shear and zoom introduce slightly more disruption, with cosine similarity values dropping to 0.91–0.94 for ImageNet and DIV2K. Among geometric distortions, zoom exhibits the most pronounced impact on feature consistency, likely due to scale-dependent feature extraction challenges.
@@
Table 3: Feature cosine similarity and image reconstruction PSNR under geometric transformations (mean $\pm$ std over 5 runs).
Dataset Cosine Similarity $\uparrow$ PSNR (dB) $\uparrow$
@@
MSCOCO $0.97\pm 0.012$ $0.96\pm 0.018$ $0.96\pm 0.017$ $0.94\pm 0.021$ $0.95\pm 0.019$ $33.5\pm 0.38$ $31.1\pm 0.56$ $31.6\pm 0.49$ $32.3\pm 0.61$ $29.6\pm 0.74$
ImageNet $0.95\pm 0.019$ $0.94\pm 0.024$ $0.94\pm 0.027$ $0.91\pm 0.033$ $0.93\pm 0.026$ $1.9\pm 0.64 33$$3.5\pm 0.51$ $29.2\pm 0.97$ $31.8\pm 0.68$ $29.1\pm 0.86$
Flickr30k $0.97\pm 0.013$ $0.95\pm 0.019$ $0.94\pm 0.028$ $0.94\pm 0.022$ $0.92\pm 0.027$ $0.2\pm 0.94 33$$3.7\pm 0.63$ $31.7\pm 1.21$ $30.2\pm 0.88$ $28.6\pm 1.09$
DIV2K $0.96\pm 0.021$ $0.92\pm 0.030$ $0.93\pm 0.027$ $0.92\pm 0.031$ $0.94\pm 0.024$ $31.5\pm 0.82$ $32.1\pm 0.71$ $30.8\pm 1.18$ $32.8\pm 0.93$ $29.6\pm 1.11$
The results show that the proposed framework effectively learns invariant features that generalize well to unseen distortions, including common image processing and geometric transformations. This robustness stems from the adversarial training setup, where the generator produces distortion-invariant features that the discriminator cannot distinguish. By aligning features across original and distorted images, the model learns to capture semantic content rather than distortion-specific cues. As a result, the features remain stable even under novel distortions at test time, enabling reliable watermark verification. To demonstrate that these features retain semantic information, reconstruction results are shown in Figure 6.
@@
4.3 Comparative Analysis of the Proposed Invariant Feature
To evaluate the robustness and effectiveness of the proposed invariant feature learning framework, we compare it with five representative self-supervised learning (SSL) baselines: SimCLR [4], BYOL [11], Barlow Twins [27], VICReg [2], and VIBCREG [17]. These methods were selected for their strong performance across SSL benchmarks and for covering diverse learning paradigms, including contrastive learning, redundancy reduction, and variance-invariance-covariance regularization. We assess the learned representations along two key axes: (1) feature invariance, i.e., the stability of feature embeddings under distortions, and (2) semantic descriptiveness, i.e., how well the features support classification when linearly probed. These two criteria together offer a comprehensive measure of robustness and utility for downstream watermarking tasks.
4.3.1 Invariant Feature Similarity
@@
Although self-supervised learning (SSL) models provide strong general-purpose visual representations, their invariance is primarily driven by predefined augmentations and does not explicitly target distortion-specific variability that affects watermark consistency. As a result, SSL features may still exhibit instability under distortions such as compression, filtering, and geometric transformations.
In contrast, the proposed framework explicitly enforces distortion invariance through adversarial training, which minimizes feature discrepancies between clean and distorted inputs. At the same time, the reconstruction constraint ensures that the learned features retain sufficient semantic information and avoid collapse. This targeted invariance learning leads to more stable and reliable feature representations for multi-bit zero-watermark decoding under diverse distortions.
Table 4 reports the results across four datasets. The proposed method consistently outperforms the SSL baselines across all distortions and datasets. For instance, under Gaussian blur and JPEG compression, our model achieves cosine similarities up to 0.98, demonstrating high resistance to low-level degradations. Even under challenging conditions like salt & pepper noise and hue shifts, our model maintains higher feature consistency than all compared methods. These results validate the effectiveness of our adversarial invariant learning objective in producing distortion-invariant feature embeddings.
Table 4: Quantitative Comparison of Feature Invariance (Cosine Similarity) across SSL Methods Under Diverse Distortions on Multiple Datasets.
@@
4.3.2 Linear Evaluation for Semantic Discriminability
@@
4.4 Robust Zero Watermarking
This subsection presents an evaluation of the proposed learning-based zero-watermarking scheme using the invariant features extracted in Module 1. We organize the analysis into three parts. First, we visualize the training dynamics of the reference signature $C$, including t-SNE projections before and after optimization, and monitor the convergence behavior of the watermark loss over time. Next, we quantify the robustness of the watermarking system across a wide range of image distortions by evaluating bit accuracy under varying transformation intensities. Finally, we compare our method against existing state-of-the-art zero-watermarking and other deep watermarking approaches, highlighting the performance gains achieved by our framework in terms of robustness and stability. While most previous studies have relied on pretrained models to extract features for robust zero-watermarking, we used our own invariant feature model to enhance robustness. Together, these experiments establish the practicality and generalizability of our proposed method. We evaluated the proposed method using binary watermark messages of varying lengths and report results using a 30-bit watermark to align with common practice in the literature and ensure fair comparison with prior work.
4.4.1 Visualization of Reference Signature Learning
@@
As shown in Fig. 9, we apply t-distributed Stochastic Neighbor Embedding (t-SNE) to project the high-dimensional reference vectors $C\in\mathbb{R}^{k\times d}$ into two dimensions, both before and after training. Each point represents the projection of one bit vector $C_{i}$ associated with a distorted image. Before training, the projected vectors appear scattered without clear structure, reflecting the random initialization. After optimization, the vectors form distinct and compact clusters in the embedding space. This emergent separation indicates that the learning process successfully aligns the reference signature with the distortion-invariant feature space. Notably, the improved spatial organization mirrors the expected behavior of our theoretical formulation, in which $C$ is optimized to serve as a directional codebook that maximizes the margin between binary watermark bits. The optimization implicitly identifies a hyperplane such that $\tilde{F}\cdot C_{i}\geq 0.5$ or <0.5<0.5, depending on the bit label, enabling robust watermark recovery through simple dot-product thresholding.
@@
4.4.2 Distortion Tolerance Range
@@
For JPEG compression, BER consistently decreases as the quality factor improves, confirming strong robustness to compression artifacts. Gaussian blur exhibits a gradual BER increase with higher blur radii, while cropping and hue adjustment maintain relatively low BER values even at high distortion intensities. This indicates that the proposed method effectively preserves watermark integrity under spatial and color-shift perturbations. Noise-based distortions, especially Gaussian noise and salt & pepper noise, introduce relatively more significant degradation, with BER rising sharply under higher noise levels. Nonetheless, the absolute BER values remain low across all cases, ranging from 0.05% to a maximum of 8%, highlighting the overall robustness of the proposed method.
4.4.3 Comparison with Existing Watermarking Methods
Table 5 highlights the key distinctions between the proposed framework and existing approaches. Traditional and deep zero-watermarking methods rely on fixed or task-driven features without explicitly enforcing distortion invariance, resulting in limited robustness under diverse corruptions. SSL-based representations provide general semantic invariance but are not optimized to suppress distortion-specific variability that directly affects watermark signatures.
Table 5: Comparison of watermarking paradigms and invariant representation properties.
Method Training Paradigm Invariance Mechanism Feature Stability Generalization to Unseen Distortions Multi-bit
Traditional Zero-WM [1] Handcrafted None (fixed design) Limited, distortion-specific Weak Limited
Deep Zero-WM [10] Task-driven CNN Implicit (data-dependent) Moderate Limited
SSL Feature-based [7] Pre-trained (task-agnostic) Augmentation invariance Semantic but not distortion-stable Moderate Limited
Deep WM (Embedding) [31] End-to-end embedding Learned robustness (task-specific) In-domain Limited (overfits distortions)
InvZW (Ours) Invariant representation learning Adversarial + reconstruction High, distortion-invariant Strong (unseen distortions)
In contrast, the proposed framework explicitly learns distortion-invariant representations through adversarial alignment between clean and distorted inputs, while a reconstruction constraint preserves semantic structure and prevents feature collapse. This combination enables stable feature representations that generalize beyond the distortions observed during training, providing improved robustness for multi-bit zero-watermarking under diverse real-world conditions.
To quantitatively evaluate the robustness of the proposed watermarking framework, we compare it against several representative models across three categories: (1) deep learning-based watermarking schemes that embed information into the image, (2) feature extractor-based watermarking methods that leverage pre-trained SSL models, and (3) zero-watermarking approaches with deep models. Each comparison focuses on bit-level accuracy or BER under various common distortions.
For a fair comparison, all methods were evaluated using a fixed watermark length of 30 bits per image. Additionally, watermarking embedding was controlled to ensure consistency across methods. In particular, for embedded watermarking schemes, watermarking were applied such that the resulting marked images maintained a PSNR of approximately 40 dB. This constraint was applied to ensure consistent visual quality across all methods, enabling a fair evaluation of robustness under similar imperceptibility conditions. Although some existing methods do not adopt this practice, using their original configurations could result in marked images with varying PSNR levels, leading to an imbalanced comparison. By standardizing the distortion budget, we provide a more objective and controlled benchmarking environment across both embedded and zero-watermarking schemes. In contrast, the proposed method and other zero-watermarking baselines do not alter the original image, preserving its fidelity by design.
Table 6: Bit Accuracy (%) Comparison with Watermarking Models that Embed Information.
@@ 4.4.3 Comparison with Existing Watermarking Methods
Table 6 presents bit accuracy results under a wide range of distortions. We compare our method with HIDDEN [31], a widely used encoder-decoder watermarking scheme, and [20], which uses adversarial training to improve robustness. Our model consistently achieves the highest accuracy across most transformations, particularly under color-related distortions such as hue and saturation adjustment, where alternative methods degrade significantly. Moreover, our model shows stronger robustness under severe noise (e.g., Gaussian 0.10, Salt & Pepper 0.15), outperforming others by large margins. The average accuracy improvement of our model demonstrates superior generalization across both geometric and photometric distortions.
Table 7: Bit Accuracy (%) Comparison with Pretrained Feature-Based Watermarking Models.
@@
Table 7 compares our framework with two recent methods that utilize pre-trained SSL models for watermark retrieval: Vukotić et al. [24], which uses convolutional networks, and Fernández et al. [7], which employs DINO features. Although these approaches demonstrate strong performance under identity-preserving transformations, they often struggle with heavy attacks and geometric alterations such as rotation or crop. In contrast, our model maintains perfect or near-perfect bit accuracy across all comparison distortions. Notably, under non-differentiable transformations like screenshots, our method preserves 98% bit accuracy, further validating the resilience conferred by our invariant feature learning.
Table 8: BER (%) Comparison across Zero-Watermarking Models.
@@
JPEG Compression (30) 0.0267 0.0431
JPEG Compression (70) 0.0169 0.0748
@@
Table 8 reports BER comparisons with a recent deep zero-watermarking method from [13]. Across several distortion types, including geometric transformations and Gaussian filtering, our method demonstrates competitive robustness and often achieves lower BER than the baseline. In particular, under rotation and scaling, as well as Gaussian filtering, the proposed framework consistently yields improved performance.
At the same time, the baseline method performs better under certain distortions, notably JPEG compression and median filtering. In particular, under strong JPEG compression (e.g., quality 30 and 70), the proposed method exhibits higher BER, indicating that although compression artifacts are captured by the learned invariant features, they are not explicitly optimized in the same way as handcrafted or distortion-specific designs. These differences reflect the distinct design principles of the two approaches. Existing methods rely on fixed or handcrafted feature stability, which can be effective under specific distortions such as compression, while our method focuses on learning invariant representations that generalize across diverse transformations, particularly geometric and filtering distortions.
Overall, the results suggest that the proposed invariant feature framework provides a flexible and generalizable representation, achieving strong performance across a wide range of distortions while exhibiting complementary behavior to existing methods.
4.5 Ablation Study
To assess the importance of the Reconstructor module $R$, we conduct an ablation study by removing $R$ from the training pipeline. As discussed in Section 3, $R$ is introduced to prevent feature collapse where the feature extractor learns to produce constant or near-zero outputs that are trivially invariant but devoid of semantic content. Although such collapsed representations may fool the distortion discriminator, they lack the richness necessary to support watermarking or downstream tasks. The reconstructor acts as a semantic regularizer, ensuring that invariant features must still encode sufficient information to reconstruct the input image. This auxiliary constraint discourages degenerate solutions by penalizing high reconstruction error when semantic content is not preserved.
Table 9 shows a substantial drop in both linear classification accuracy (obtained by training a shallow classifier on top of frozen features) and watermark bit accuracy when the Reconstructor is removed. Specifically, linear accuracy drops from 81.67% to 63.31%, and bit accuracy degrades from 96.7% to 68.1%. These declines indicate that the learned features lose both discriminative and semantic structure in the absence of $R$, which directly impacts its ability to support watermark extraction or classification.
Table 9: Reconstructor Improves Feature Semantics and Watermark Recovery
Reconstructor Linear Accuracy (↑) Bit Accuracy (↑)
$\times$ 63.31% 68.1%
81.67% 96.7%
To assess feature diversity, we compute the average pairwise cosine similarity between features extracted from different input images across the testing sets. Fig. 12 shows heatmaps for models trained with and without the Reconstructor. With the Reconstructor, features exhibit strong diagonal patterns and varied similarities, indicating semantic distinction across inputs. Without it, features become nearly identical, resulting in uniformly high similarity - clear evidence of feature collapse.
Figure 12: Cosine similarity heatmaps for 30 input images. Left: With Reconstructor, feature vectors are diverse and input-dependent. Right: Without Reconstructor, features collapse to nearly identical outputs.
Table 10: Adversarial Training Improves Feature Invariance and Watermark Recovery
Adversarial Linear Accuracy (↑) Bit Accuracy (↑)
$\times$ 69.45% 74.3%
81.67% 96.7%
To evaluate the role of adversarial training, we compare models trained with and without the adversarial objective while keeping other components unchanged. As shown in Table 10, removing adversarial training leads to a noticeable drop in both linear evaluation accuracy and watermark bit accuracy. This indicates that, without adversarial supervision, the learned features are less robust to distortions and less effective for reliable watermark decoding. In contrast, incorporating adversarial training significantly improves both semantic representation quality and watermark recovery performance, demonstrating its critical role in enforcing distortion-invariant features, preventing feature collapse and ensuring meaningful invariant representations.
In addition, these findings empirically validate the theoretical formulation in Eq. 5, where the goal is to maximize the mutual information $I(F,x)$ between the learned features $F$ and the original image $x$, while minimizing $I(F,x^{\prime})$ with respect to distorted inputs $x^{\prime}$. The reconstruction loss enforces the first objective by requiring that $F$ retain enough signal to recover $x$, thereby serving as a proxy for maximizing $I(F,x)$. Without $R$, this semantic pathway is lost, and the model is free to collapse to trivial invariants. Thus, the reconstructor plays an essential role in balancing invariance and informativeness, acting as a critical regularizer for robust and meaningful representation.
4.6 Computational Cost and Scalability
The proposed framework consists of a ViT-based feature extractor, a discriminator, and a reconstructor. The feature extractor contains approximately 9M parameters, while the discriminator and reconstructor contain approximately 14M and 7M parameters, respectively. Despite the multi-component design, all modules operate on relatively small inputs of size $128\times 128$, making training computationally practical. In our implementation, the full training process is conducted on a single GPU (Nvidia RTX 4090) and takes approximately 46 hours. The framework can be further scaled by replacing the backbone with larger pre-trained models, without modifying the overall pipeline.
At inference time, only the feature extractor is required for watermark extraction, while the discriminator and reconstructor are used during training only. This significantly reduces the deployment cost and allows efficient inference comparable to standard feature extraction pipelines.
In terms of storage, the framework requires storing a reference signature tensor of size $128\times 128\times 3$ for each image–watermark pair. The storage cost is therefore linear in the number of protected samples and depends on the numerical precision used for the stored features.
5 Conclusion